Calculate the length of a constant wire of cross sectional areas 4π x 10^8m and resistivity 1.1 x 10^-6Ωm required to construct a standard resistor of resistance 21Ω (take π=22/7)
1
Expert's answer
2020-05-27T10:32:47-0400

Resistance of wire of length "l", cross sectional area "S" and resistivity "\rho" is "R = \frac{\rho l}{S}", from where "l = \frac{ R S}{\rho} = \frac{21 \Omega \cdot 4\pi \cdot 10^{-8}m^2}{1.1 \cdot 10^{-6} \Omega \cdot m} = 2.4 m".